OpenStudy (anonymous):

Long Division Help! I forgot how to long divide and would like to learn. So how would you long divide 25378/184?

OpenStudy (anonymous):

Please show work and explain steps

OpenStudy (anonymous):

https://www.youtube.com/watch?v=eIUoIhfupuA

OpenStudy (amistre64):

one what is to x=10 ----------------------------- x^2+8x +4 | 2x^4 + 5x^3 + 3x^2 + 7x + 8

OpenStudy (anonymous):

@amistre64 I do not understand that

OpenStudy (amistre64):

then you might have to just work it like we did back in the 3rd or 4th grade ....

OpenStudy (anonymous):

Yeah that is actually what I wanted, a refresh mostly

OpenStudy (amistre64):

hard to recall which grade it was really, been sooo long ago

OpenStudy (anonymous):

But I would like to learn what you did there, I can divide polynomials using a box/grid method. Long division is really need though

OpenStudy (amistre64):

well, we find an integer k such that 253 = 184k + r for some remainder r

OpenStudy (amistre64):

if you can divide polys, divide the poly i presented

OpenStudy (anonymous):

Okay, but I do now know how to do it with long division, the grid isn't accepted to my teacher

OpenStudy (amistre64):

find an integer k such that 253 = 184k + r for some remainder r in other words we want 184k to be just smaller than 253

OpenStudy (anonymous):

k=1 and r=69?

OpenStudy (amistre64):

1 is good, 253 -184 ----- 69 thats good too, now drop down the next digit and do it again

OpenStudy (amistre64):

1 -------- 184 | 25378 (184) ------- 697 697 = 184k + r for some other k and r this time

OpenStudy (anonymous):

So 697=184k+r k=3 and r=145?

OpenStudy (anonymous):

Does k have to be a whole number?

OpenStudy (amistre64):

not really, but the algorithm is simpler if we let it be a whole number

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So where are we at in terms of the 

OpenStudy (amistre64):

for example 127 = 1(10^2) + 2(10^1) + 7(10)^0 but it could just as well be 127 = 2/3(10^2) + 1/5(10^1) + 175/3(10)^0

OpenStudy (amistre64):

13 -------- 184 | 25378 (184)vv ----- 697v (552)v ---- 1458

OpenStudy (anonymous):

This time k=7 r=170

OpenStudy (anonymous):

So just repeat and add decimal later?

OpenStudy (amistre64):

then we have come to the end of it 170 out of 184 is left over

OpenStudy (amistre64):

if youwant to decimate it sure ....

OpenStudy (anonymous):

Okay great, now what were you doing with the polys and how did you make them?

OpenStudy (amistre64):

25378 = 184(137) + 170

OpenStudy (amistre64):

well since our number system is base 10 each position has a digit that is multiplied by a power of 10 instead of 10, let it be x and you have a polynomial representaiton of the numbers

OpenStudy (anonymous):

Oh okay that is easy, then divide the polys, and what does the answer of the polys mean? Same base 10 concept?

OpenStudy (amistre64):

yeah, itll be the same number but in a base x representation, let x = 10 again 2x^2 -11x +83 ----------------------------- x^2+8x +4 | 2x^4 + 5x^3 + 3x^2 + 7x + 8 (2x^4+16x^3 +8x^2) -------------------- -11x^3 -5x^2 +7x +8 (-11x^3-88x^2 -44x) -------------------- 83x^2 + 51x +8 (83x^2+664x +332) ------------------ -613x -324 200 - 110 + 83 = 173 not too sure if i got the last part (the remainder) worked out right tho

OpenStudy (anonymous):

Oh okay, thank you so very much for the great amount of help! Okay and final question, I know I won't be any time soon, but how does one become a moderator?

OpenStudy (amistre64):

one does not become a moderator, one is chosen for the position.

OpenStudy (anonymous):

Oh okay, I thanks for clearing that up. Okay so thank you very much for the help!

OpenStudy (amistre64):

youre welcome